OpenStudy (anonymous):

Solve for x: -2|x-3|=-12 I need help somebody!!

OpenStudy (akashdeepdeb):

-2 |x-3| = -12 2 |x-3| = 12 |x-3| = 6 Now what values do you think x can take? :)

OpenStudy (anonymous):

Idk... I have a hard time with math :/ I feel like an idiot.

OpenStudy (akashdeepdeb):

Well, then we are in this together because I feel like one too, a different type though. Do you know what, |x| means? Or what do these: | | symbols represent?